# Hypothesis Spaces for Deep Learning

### 05 Mar 2024 v1

@@ 2 Learning with Deep Neural Networks
For $k$ vector-valued functions $f_{j}$, $j\in\mathbb{N}_{k}$, where the range of $f_{j}$ is contained in the domain of $f_{j+1}$, for $j\in\mathbb{N}_{k-1}$, we denote the consecutive composition of $f_{j}$, $j\in\mathbb{N}_{k}$, by
@@ 3 Vector-Valued Reproducing Kernel Banach Space
It is worth of pointing out that although $\mathcal{B}$ is a space of vector-valued functions, the $\delta$-dual space $\mathcal{B}^{\prime}$ defined here is a space of scalar-valued functions. This is determined by the form of the point evaluation functionals in set $\Delta$ defined by (13). The way of defining the dual space of the vector-valued RKBS $\mathcal{B}$ is not unique. One can also define a dual space of the vector-valued RKBS $\mathcal{B}$ as a space of vector-valued functions. In this paper, we adopt the current form of $\mathcal{B}^{\prime}$ since it is simple and sufficient to serve our purposes. Other forms of the $\delta$-dual space will be investigated in a different occasion.
@@ 4 Hypothesis Space
is an isometric isomorphism. As has been shown earlier, the map $\varphi$ defined by (25) is a isometric isomorphism from $\mathcal{B}_{\mathcal{N}}$ to $\mathcal{M}(\Theta)/\mathcal{S}^{\perp}$. As a result, $\tau\circ\varphi$ provides an isometric isomorphism from $\mathcal{B}_{\mathcal{N}}$ to $\mathcal{S}^{*}$. ∎
@@ 5 Representer Theorems for Learning Solutions
Clearly, the MNI problem (45) include $tm$ interpolation conditions which are produced by the linear functionals in the set
@@
then for any extreme point $\hat{f}$ of the solution set of the MNI problem (46), there exist $\gamma_{j}\geq 0$, $j\in\mathbb{N}_{n}$, with $\sum_{j\in\mathbb{N}_{n}}\gamma_{j}=\|\hat{\nu}\|_{B_{*}}$ and $u_{j}\in\mathrm{ext}\left(\partial\|\cdot\|_{B_{*}}(\hat{\nu})\right)$, $j\in\mathbb{N}_{n}$, such that
@@
Suppose that $m$ distinct points $x_{j}\in\mathbb{R}^{s}$, $j\in\mathbb{N}_{m}$, and $\mathbf{Y}\in\mathbb{R}^{t\times m}\backslash\{\mathbf{0}\}$ are given, and the functionals in $\mathbb{K}_{\mathcal{X}}$ are linearly independent. Let $\hat{g}$ be the function defined by (50). Then for any $\hat{f}\in\mathrm{ext}(\mathbb{S}_{\mathcal{X},\mathbf{Y}})$, there exist $\gamma_{\ell}\geq 0$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $h_{\ell}\in\mathrm{ext}(\partial\|\cdot\|_{\infty}(\hat{g}))$, $\ell\in\mathbb{N}_{tm}$, such that
@@
Hence, the hypothesis of Lemma 12 is satisfied. Then by Lemma 12, we can represent any extreme point $\hat{f}$ of the solution set $\mathbb{S}_{\mathcal{X},\mathbf{Y}}$ of the MNI problem (45) as in equation (51) for some $\gamma_{\ell}\geq 0$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $h_{\ell}\in\mathrm{ext}(\partial\|\cdot\|_{\infty}(\hat{g}))$, $\ell\in\mathbb{N}_{tm}$. ∎
@@
Suppose that $m$ distinct points $x_{j}\in\mathbb{R}^{s}$, $j\in\mathbb{N}_{m}$, and $\mathbf{Y}\in\mathbb{R}^{t\times m}\backslash\{\mathbf{0}\}$ are given, and the functionals in $\mathbb{K}_{\mathcal{X}}$ are linearly independent functions in $C_{0}(\Theta)$. If ${\hat{\mu}}$ is a solution of the MNI problem (54), then $f_{\hat{\mu}}(x):=\left[f_{\hat{\mu}}^{k}(x):k\in\mathbb{N}_{t}\right]^{\top}$, $x\in\mathbb{R}^{s}$, with $f_{\hat{\mu}}^{k}$, $k\in\mathbb{N}_{t}$, defined as in (20) with $\mu$ replaced by $\hat{\mu}$, is a solution of the MNI problem (45) and $\|f_{{\hat{\mu}}}\|_{\mathcal{B}_{\mathcal{N}}}=\|{\hat{\mu}}\|_{\mathrm{TV}}$.
@@
Suppose that $m$ distinct points $x_{j}\in\mathbb{R}^{s}$, $j\in\mathbb{N}_{m}$, and $\mathbf{Y}\in\mathbb{R}^{t\times m}\backslash\{\mathbf{0}\}$ are given, and the functionals in $\mathbb{K}_{\mathcal{X}}$ are linearly independent. Let $\hat{g}$ be the function defined by (50). Then for any $\hat{\mu}\in\mathrm{ext}\left(\widetilde{\mathbb{S}}_{\mathcal{X},\mathbf{Y}}\right)$, there exist $\gamma_{\ell}\geq 0$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $\theta_{\ell}\in\Theta(\hat{g})$, $\ell\in\mathbb{N}_{tm}$, such that
@@
in which the subdifferential set $\partial\|\cdot\|_{\infty}(\hat{g})$ is a subset of the measure space $\mathcal{M}(\Theta)$. As a result, the hypothesis of Lemma 12 is satisfied. According to Lemma 12, for any $\hat{\mu}\in\mathrm{ext}\left(\widetilde{\mathbb{S}}_{\mathcal{X},\mathbf{Y}}\right)$, there exist $\gamma_{\ell}\geq 0$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $u_{\ell}\in\mathrm{ext}(\partial\|\cdot\|_{\infty}(\hat{g}))$, $\ell\in\mathbb{N}_{tm}$, such that
@@
for some $\gamma_{\ell}\geq 0$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $\theta_{\ell}\in\Theta(\hat{g})$, $\ell\in\mathbb{N}_{tm}$.
@@
By Proposition 1 of [23], the MNI problem (54) has at least one solution. That is, $\widetilde{\mathbb{S}}_{\mathcal{X},\mathbf{Y}}$ is nonempty and moreover, $\mathrm{ext}\left(\widetilde{\mathbb{S}}_{\mathcal{X},\mathbf{Y}}\right)$ is nonempty. We choose $\hat{\mu}\in\mathrm{ext}\left(\widetilde{\mathbb{S}}_{\mathcal{X},\mathbf{Y}}\right)$. Proposition 15 ensures that there exist $\gamma_{\ell}\geq 0$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $\theta_{\ell}\in\Theta(\hat{g})$, $\ell\in\mathbb{N}_{tm}$, such that $\hat{\mu}$ may be expressed as in equation (62). Since $\hat{\mu}$ is a solution of problem (54), we get by Proposition 14 that $f_{\hat{\mu}}=[f_{\hat{\mu}}^{k}:k\in\mathbb{N}_{t}]$ is a solution of the MNI problem (45). By definition (20) of $f_{\hat{\mu}}^{k}$, $k\in\mathbb{N}_{t}$, we have that
@@
    for some $\hat{g}\in\mathcal{V}_{\mathcal{N}}$, $\gamma_{\ell}\geq 0$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $h_{\ell}\in\mathrm{ext}(\partial\|\cdot\|_{\infty}(\hat{g}))$, $\ell\in\mathbb{N}_{tm}$.
@@
    If the loss function $\mathcal{Q}$ is convex, then every nonzero extreme point $\hat{f}$ of $\mathbb{R}_{\mathcal{X},\mathbf{Y}}$ has the form of (69) for some $\hat{g}\in\mathcal{V}_{\mathcal{N}}$, $\gamma_{\ell}\geq 0$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $h_{\ell}\in\mathrm{ext}(\partial\|\cdot\|_{\infty}(\hat{g}))$, $\ell\in\mathbb{N}_{tm}$.
@@
We first prove Item 1. Note that $\mathbb{R}_{\mathcal{X},\mathbf{Y}}$ is an nonempty set. It follows from the hypothesis $\mathcal{D}_{\mathcal{X},\mathbf{Y}}\neq\{\mathbf{0}\}$ that there exists $\hat{h}\in\mathbb{R}_{\mathcal{X},\mathbf{Y}}$ such that $\hat{\mathbf{Z}}:=\mathbf{I}_{\mathcal{X}}(\hat{h})\neq 0$. According to equation (67) with noting that $\hat{\mathbf{Z}}\in\mathcal{D}_{\mathcal{X},\mathbf{Y}}$, we have that $\mathbb{S}_{\mathcal{X},\hat{\mathbf{Z}}}\subset\mathbb{R}_{\mathcal{X},\mathbf{Y}}$ and thus, $\mathrm{ext}\left(\mathbb{S}_{\mathcal{X},\hat{\mathbf{Z}}}\right)\subset\mathbb{R}_{\mathcal{X},\mathbf{Y}}$. We choose $\hat{f}\in\mathrm{ext}\left(\mathbb{S}_{\mathcal{X},\hat{\mathbf{Z}}}\right)$ and verify that $\hat{f}$ can be represented as in (69). To this end, we choose $\hat{\mathbf{c}}:=[\hat{c}_{kj}:k\in\mathbb{N}_{t},j\in\mathbb{N}_{m}]\in\mathbb{R}^{t\times m}$ to be a solution of the dual problem (49) with $\mathbf{Y}$ replaced by $\hat{\mathbf{Z}}$. Let $\hat{g}$ be the function defined by (50) with $\hat{\mathbf{c}}$. Theorem 13 ensures that $\hat{f}$, as an extreme point of the solution set $\mathbb{S}_{\mathcal{X},\hat{\mathbf{Z}}}$, can be represented as in (69) for some $\gamma_{\ell}\geq 0$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $h_{\ell}\in\mathrm{ext}(\partial\|\cdot\|_{\infty}(\hat{g}))$, $\ell\in\mathbb{N}_{tm}$.
We next show Item 2. Assume that $\hat{f}$ is an arbitrary nonzero extreme point of $\mathbb{R}_{\mathcal{X},\mathbf{Y}}$, that is $\hat{f}\in\mathrm{ext}(\mathbb{R}_{\mathcal{X},\mathbf{Y}})\backslash\{0\}$. Because the loss function $\mathcal{Q}$ is convex, the inclusion relation (68) is satisfied. By (68), there exists $\hat{\mathbf{Z}}\in\mathcal{D}_{\mathcal{X},\mathbf{Y}}$ such that $\hat{f}\in\mathrm{ext}\left(\mathbb{S}_{\mathcal{X},\hat{\mathbf{Z}}}\right)$. Clearly, $\hat{\mathbf{Z}}\neq 0.$ Assume to the contrary that $\hat{\mathbf{Z}}=0.$ We then must have that $\mathbb{S}_{\mathcal{X},\hat{\mathbf{Z}}}=\{0\}.$ As a result, $\hat{f}=0,$ which is a contradiction. Again, let $\hat{g}$ be defined by (50) with $\hat{\mathbf{c}}$ being a solution of problem (49) with $\mathbf{Y}$ being replaced by $\hat{\mathbf{Z}}$. By Theorem 13, we can represent $\hat{f}$ as in (69) for some $\gamma_{\ell}\geq 0$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $h_{\ell}\in\mathrm{ext}(\partial\|\cdot\|_{\infty}(\hat{g}))$, $\ell\in\mathbb{N}_{tm}$. ∎
@@
for some $\hat{g}\in\mathcal{V}_{\mathcal{N}}$, $\gamma_{\ell}\geq 0$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $\theta_{\ell}\in\Theta(\hat{g})$, $\ell\in\mathbb{N}_{tm}$,
@@
Since $\mathbb{R}_{\mathcal{X},\mathbf{Y}}$ is nonempty and $\mathcal{D}_{\lambda,\mathbf{Y}}\neq\{\mathbf{0}\}$, there exists $\hat{h}\in\mathbb{R}_{\mathcal{X},\mathbf{Y}}$ such that $\hat{\mathbf{Z}}:=\mathbf{I}_{\mathcal{X}}(\hat{h})\neq\mathbf{0}$. We choose $\hat{g}$ in the form of (50), where $\hat{\mathbf{c}}$ is a solution of problem (49) with $\mathbf{Y}$ being replaced by $\hat{\mathbf{Z}}$. According to Theorem 16, the MNI problem (45) with $\mathbf{Y}:=\hat{\mathbf{Z}}$ has a solution $\hat{f}$ in the form of (70), for some $\gamma_{\ell}\geq 0$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $\theta_{\ell}\in\Theta(\hat{g})$, $\ell\in\mathbb{N}_{tm}$. In other words, $\hat{f}\in\mathbb{S}_{\mathcal{X},\hat{\mathbf{Z}}}$ and it has the form of (70). It follows from relation (67) that $\mathbb{S}_{\mathcal{X},\hat{\mathbf{Z}}}\subset\mathbb{R}_{\mathcal{X},\mathbf{Y}}$, which implies that $\hat{f}\in\mathbb{R}_{\mathcal{X},\mathbf{Y}}$. ∎

### 11 Mar 2024 v2

@@ 2 Learning with Deep Neural Networks
For each $n\in\mathbb{N}$, let $\mathbb{N}_{n}:=\{1,2,\ldots,n\}$. For $k$ vector-valued functions $f_{j}$, $j\in\mathbb{N}_{k}$, where the range of $f_{j}$ is contained in the domain of $f_{j+1}$, for $j\in\mathbb{N}_{k-1}$, we denote the consecutive composition of $f_{j}$, $j\in\mathbb{N}_{k}$, by
@@ 3 Vector-Valued Reproducing Kernel Banach Space
It is worth of pointing out that although $\mathcal{B}$ is a space of vector-valued functions, the $\delta$-dual space $\mathcal{B}^{\prime}$ defined here is a space of scalar-valued functions. This is determined by the form of the point evaluation functionals in set $\Delta$ defined by (13). The way of defining the $\delta$-dual space of the vector-valued RKBS $\mathcal{B}$ is not unique. One can also define a $\delta$-dual space of the vector-valued RKBS $\mathcal{B}$ as a space of vector-valued functions. In this paper, we adopt the current form of $\mathcal{B}^{\prime}$ since it is simple and sufficient to serve our purposes. Other forms of the $\delta$-dual space will be investigated in a different occasion.
@@ 4 Hypothesis Space
is an isometric isomorphism. As has been shown earlier, the map $\varphi$ defined by (25) is an isometric isomorphism from $\mathcal{B}_{\mathcal{N}}$ to $\mathcal{M}(\Theta)/\mathcal{S}^{\perp}$. As a result, $\tau\circ\varphi$ provides an isometric isomorphism from $\mathcal{B}_{\mathcal{N}}$ to $\mathcal{S}^{*}$. ∎
@@ 5 Representer Theorems for Learning Solutions
Clearly, the MNI problem (45) includes $tm$ interpolation conditions which are produced by the linear functionals in the set
@@
then for any extreme point $\hat{f}$ of the solution set of the MNI problem (46), there exist $\gamma_{j}\in\mathbb{R}$, $j\in\mathbb{N}_{n}$, with $\sum_{j\in\mathbb{N}_{n}}\gamma_{j}=\|\hat{\nu}\|_{B_{*}}$ and $u_{j}\in\mathrm{ext}\left(\partial\|\cdot\|_{B_{*}}(\hat{\nu})\right)$, $j\in\mathbb{N}_{n}$, such that
@@
Suppose that $m$ distinct points $x_{j}\in\mathbb{R}^{s}$, $j\in\mathbb{N}_{m}$, and $\mathbf{Y}\in\mathbb{R}^{t\times m}\backslash\{\mathbf{0}\}$ are given, and the functionals in $\mathbb{K}_{\mathcal{X}}$ are linearly independent. Let $\hat{g}$ be the function defined by (50). Then for any $\hat{f}\in\mathrm{ext}(\mathbb{S}_{\mathcal{X},\mathbf{Y}})$, there exist $\gamma_{\ell}\in\mathbb{R}$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $h_{\ell}\in\mathrm{ext}(\partial\|\cdot\|_{\infty}(\hat{g}))$, $\ell\in\mathbb{N}_{tm}$, such that
@@
Hence, the hypothesis of Lemma 12 is satisfied. Then by Lemma 12, we can represent any extreme point $\hat{f}$ of the solution set $\mathbb{S}_{\mathcal{X},\mathbf{Y}}$ of the MNI problem (45) as in equation (51) for some $\gamma_{\ell}\in\mathbb{R}$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $h_{\ell}\in\mathrm{ext}(\partial\|\cdot\|_{\infty}(\hat{g}))$, $\ell\in\mathbb{N}_{tm}$. ∎
@@
Suppose that $m$ distinct points $x_{j}\in\mathbb{R}^{s}$, $j\in\mathbb{N}_{m}$, and $\mathbf{Y}\in\mathbb{R}^{t\times m}\backslash\{\mathbf{0}\}$ are given, and the functionals in $\mathbb{K}_{\mathcal{X}}$ are linearly independent. If ${\hat{\mu}}$ is a solution of the MNI problem (54), then $f_{\hat{\mu}}(x):=\left[f_{\hat{\mu}}^{k}(x):k\in\mathbb{N}_{t}\right]^{\top}$, $x\in\mathbb{R}^{s}$, with $f_{\hat{\mu}}^{k}$, $k\in\mathbb{N}_{t}$, defined as in (20) with $\mu$ replaced by $\hat{\mu}$, is a solution of the MNI problem (45) and $\|f_{{\hat{\mu}}}\|_{\mathcal{B}_{\mathcal{N}}}=\|{\hat{\mu}}\|_{\mathrm{TV}}$.
@@
Suppose that $m$ distinct points $x_{j}\in\mathbb{R}^{s}$, $j\in\mathbb{N}_{m}$, and $\mathbf{Y}\in\mathbb{R}^{t\times m}\backslash\{\mathbf{0}\}$ are given, and the functionals in $\mathbb{K}_{\mathcal{X}}$ are linearly independent. Let $\hat{g}$ be the function defined by (50). Then for any $\hat{\mu}\in\mathrm{ext}\left(\widetilde{\mathbb{S}}_{\mathcal{X},\mathbf{Y}}\right)$, there exist $\gamma_{\ell}\in\mathbb{R}$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $\theta_{\ell}\in\Theta(\hat{g})$, $\ell\in\mathbb{N}_{tm}$, such that
@@
in which the subdifferential set $\partial\|\cdot\|_{\infty}(\hat{g})$ is a subset of the measure space $\mathcal{M}(\Theta)$. As a result, the hypothesis of Lemma 12 is satisfied. According to Lemma 12, for any $\hat{\mu}\in\mathrm{ext}\left(\widetilde{\mathbb{S}}_{\mathcal{X},\mathbf{Y}}\right)$, there exist $\gamma_{\ell}\in\mathbb{R}$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $u_{\ell}\in\mathrm{ext}(\partial\|\cdot\|_{\infty}(\hat{g}))$, $\ell\in\mathbb{N}_{tm}$, such that
@@
for some $\gamma_{\ell}\in\mathbb{R}$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $\theta_{\ell}\in\Theta(\hat{g})$, $\ell\in\mathbb{N}_{tm}$.
@@
By Proposition 1 of [23], the MNI problem (54) has at least one solution. That is, $\widetilde{\mathbb{S}}_{\mathcal{X},\mathbf{Y}}$ is nonempty and moreover, $\mathrm{ext}\left(\widetilde{\mathbb{S}}_{\mathcal{X},\mathbf{Y}}\right)$ is nonempty. We choose $\hat{\mu}\in\mathrm{ext}\left(\widetilde{\mathbb{S}}_{\mathcal{X},\mathbf{Y}}\right)$. Proposition 15 ensures that there exist $\gamma_{\ell}\in\mathbb{R}$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $\theta_{\ell}\in\Theta(\hat{g})$, $\ell\in\mathbb{N}_{tm}$, such that $\hat{\mu}$ may be expressed as in equation (62). Since $\hat{\mu}$ is a solution of problem (54), we get by Proposition 14 that $f_{\hat{\mu}}=[f_{\hat{\mu}}^{k}:k\in\mathbb{N}_{t}]$ is a solution of the MNI problem (45). By definition (20) of $f_{\hat{\mu}}^{k}$, $k\in\mathbb{N}_{t}$, we have that
@@
    for some $\hat{g}\in\mathcal{V}_{\mathcal{N}}$, $\gamma_{\ell}\in\mathbb{R}$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $h_{\ell}\in\mathrm{ext}(\partial\|\cdot\|_{\infty}(\hat{g}))$, $\ell\in\mathbb{N}_{tm}$.
@@
    If the loss function $\mathcal{Q}$ is convex, then every nonzero extreme point $\hat{f}$ of $\mathbb{R}_{\mathcal{X},\mathbf{Y}}$ has the form of (69) for some $\hat{g}\in\mathcal{V}_{\mathcal{N}}$, $\gamma_{\ell}\in\mathbb{R}$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $h_{\ell}\in\mathrm{ext}(\partial\|\cdot\|_{\infty}(\hat{g}))$, $\ell\in\mathbb{N}_{tm}$.
@@
We first prove Item 1. Note that $\mathbb{R}_{\mathcal{X},\mathbf{Y}}$ is an nonempty set. It follows from the hypothesis $\mathcal{D}_{\mathcal{X},\mathbf{Y}}\neq\{\mathbf{0}\}$ that there exists $\hat{h}\in\mathbb{R}_{\mathcal{X},\mathbf{Y}}$ such that $\hat{\mathbf{Z}}:=\mathbf{I}_{\mathcal{X}}(\hat{h})\neq 0$. According to equation (67) with noting that $\hat{\mathbf{Z}}\in\mathcal{D}_{\mathcal{X},\mathbf{Y}}$, we have that $\mathbb{S}_{\mathcal{X},\hat{\mathbf{Z}}}\subset\mathbb{R}_{\mathcal{X},\mathbf{Y}}$ and thus, $\mathrm{ext}\left(\mathbb{S}_{\mathcal{X},\hat{\mathbf{Z}}}\right)\subset\mathbb{R}_{\mathcal{X},\mathbf{Y}}$. We choose $\hat{f}\in\mathrm{ext}\left(\mathbb{S}_{\mathcal{X},\hat{\mathbf{Z}}}\right)$ and verify that $\hat{f}$ can be represented as in (69). To this end, we choose $\hat{\mathbf{c}}:=[\hat{c}_{kj}:k\in\mathbb{N}_{t},j\in\mathbb{N}_{m}]\in\mathbb{R}^{t\times m}$ to be a solution of the dual problem (49) with $\mathbf{Y}$ replaced by $\hat{\mathbf{Z}}$. Let $\hat{g}$ be the function defined by (50) with $\hat{\mathbf{c}}$. Theorem 13 ensures that $\hat{f}$, as an extreme point of the solution set $\mathbb{S}_{\mathcal{X},\hat{\mathbf{Z}}}$, can be represented as in (69) for some $\gamma_{\ell}\in\mathbb{R}$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $h_{\ell}\in\mathrm{ext}(\partial\|\cdot\|_{\infty}(\hat{g}))$, $\ell\in\mathbb{N}_{tm}$.
We next show Item 2. Assume that $\hat{f}$ is an arbitrary nonzero extreme point of $\mathbb{R}_{\mathcal{X},\mathbf{Y}}$, that is $\hat{f}\in\mathrm{ext}(\mathbb{R}_{\mathcal{X},\mathbf{Y}})\backslash\{0\}$. Because the loss function $\mathcal{Q}$ is convex, the inclusion relation (68) is satisfied. By (68), there exists $\hat{\mathbf{Z}}\in\mathcal{D}_{\mathcal{X},\mathbf{Y}}$ such that $\hat{f}\in\mathrm{ext}\left(\mathbb{S}_{\mathcal{X},\hat{\mathbf{Z}}}\right)$. Clearly, $\hat{\mathbf{Z}}\neq 0.$ Assume to the contrary that $\hat{\mathbf{Z}}=0.$ We then must have that $\mathbb{S}_{\mathcal{X},\hat{\mathbf{Z}}}=\{0\}.$ As a result, $\hat{f}=0,$ which is a contradiction. Again, let $\hat{g}$ be defined by (50) with $\hat{\mathbf{c}}$ being a solution of problem (49) with $\mathbf{Y}$ being replaced by $\hat{\mathbf{Z}}$. By Theorem 13, we can represent $\hat{f}$ as in (69) for some $\gamma_{\ell}\in\mathbb{R}$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $h_{\ell}\in\mathrm{ext}(\partial\|\cdot\|_{\infty}(\hat{g}))$, $\ell\in\mathbb{N}_{tm}$. ∎
@@
for some $\hat{g}\in\mathcal{V}_{\mathcal{N}}$, $\gamma_{\ell}\in\mathbb{R}$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $\theta_{\ell}\in\Theta(\hat{g})$, $\ell\in\mathbb{N}_{tm}$.
@@
Since $\mathbb{R}_{\mathcal{X},\mathbf{Y}}$ is nonempty and $\mathcal{D}_{\lambda,\mathbf{Y}}\neq\{\mathbf{0}\}$, there exists $\hat{h}\in\mathbb{R}_{\mathcal{X},\mathbf{Y}}$ such that $\hat{\mathbf{Z}}:=\mathbf{I}_{\mathcal{X}}(\hat{h})\neq\mathbf{0}$. We choose $\hat{g}$ in the form of (50), where $\hat{\mathbf{c}}$ is a solution of problem (49) with $\mathbf{Y}$ being replaced by $\hat{\mathbf{Z}}$. According to Theorem 16, the MNI problem (45) with $\mathbf{Y}:=\hat{\mathbf{Z}}$ has a solution $\hat{f}$ in the form of (70), for some $\gamma_{\ell}\in\mathbb{R}$, $\ell\in\mathbb{N}_{tm}$, with $\sum_{\ell\in\mathbb{N}_{tm}}\gamma_{\ell}=\|\hat{g}\|_{\infty}$ and $\theta_{\ell}\in\Theta(\hat{g})$, $\ell\in\mathbb{N}_{tm}$. In other words, $\hat{f}\in\mathbb{S}_{\mathcal{X},\hat{\mathbf{Z}}}$ and it has the form of (70). It follows from relation (67) that $\mathbb{S}_{\mathcal{X},\hat{\mathbf{Z}}}\subset\mathbb{R}_{\mathcal{X},\mathbf{Y}}$, which implies that $\hat{f}\in\mathbb{R}_{\mathcal{X},\mathbf{Y}}$. ∎